a history class is comprised of 8 students, 3 female, 5 male. if the instructor of the class randomly chooses 6 students from the class for an oral exam what is the probability that 2 female and 4 male students will be selected. round to 3 decimal places.
Answer #1

number of ways for two females = 3C2 =3
number of ways for four males = 5C4 =5

Total ways = 8C6 =28

Favourable ways =3+5 = 8

So P= 8/28 = 0.286